# Filtrations on derived category of twisted K3 surfaces

### v1

@@ Abstract.
In this paper, we study the Shen-Yin-Zhao type filtration on the derived category of twisted K3 surfaces. We introduce the so-called twisted Beauville-Voisin class on a twisted K3 surface and use it to extend the work in [20, 21] to the case of twisted K3 surfaces. It shares many nice properties as the untwisted case. For instance, the filtration is preserved under any Fourier-Mukai transformation. As an application, we obtain a new filtration on the Chow group of zero cycles on the Bridgeland moduli space of twisted K3 surfaces. It is birational invariant and can be viewed as a candidate for the conjectural Beauville-Voisin filtration. We compare this filtration with Voisin’s filtration and show they coincide on the locally free loci. Some other applications (such as Bloch’s conjecture for birational automorphisms) are discussed at the end.
@@ 1.1. Filtrations on K3 category
Let $X$ be a K3 surface and let $\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)$ be the Beauville-Voisin class defined in [3]. In [20], O’Grady introduced a filtration $\mathbf{S}_{\bullet}(X)$ on $\operatorname{CH}_{0}(X)$ given by
$\mathbf{S}_{i}(X)=\bigcup_{\deg([z])=i}\Big{\{}[z]+\mathbb{Z}\cdot[\mathfrak{o}_{X}]\Big{\}}$
for all effective zero-cycles $z$ of degree $i$. Denote by $\mathrm{D}^{b}(X)$ the bounded derived category of coherent sheaves on $X$. In [21], Shen-Yin-Zhao used O’Grady’s filtration to construct a natural increasing filtration on the bounded derived category $\mathrm{D}^{b}(X)$
$\mathbf{S}_{0}(\mathrm{D}^{b}(X))\subset\mathbf{S}_{1}(\mathrm{D}^{b}(X))\subset\ldots\subset\mathbf{S}_{i}(\mathrm{D}^{b}(X))\ldots$
where $\mathbf{S}_{i}(\mathrm{D}^{b}(X))$ consists of objects $\mathcal{E}\in\mathrm{D}^{b}(X)$ with $c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)$. They have shown that for an object $\mathcal{E}\in\mathrm{D}^{b}(X)$, one has
$c_{2}(\mathcal{E})\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X),$
where $d(\mathcal{E})=\frac{1}{2}\dim\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})$. More generally, let $\mathscr{X}\to X$ be a twisted K3 surface in the sense of [16]. Denote by $\mathrm{D}^{b}(\mathscr{X}):=\mathrm{D}^{(1)}(\mathscr{X})$ the bounded derived category of 1-fold twisted coherent sheaves on $\mathscr{X}\to X$ (See §2.1). It is expected that there is a similar filtration on $\mathrm{D}^{b}(\mathscr{X})$. The main result of this paper is
@@
There exists a unique cycle class $\mathfrak{o}_{\mathscr{X}}\in\operatorname{CH}_{0}(X)$ of degree $1$ such that for any $\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})$, we have
$c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X),$
where $\mathrm{d}(\mathcal{E})=\frac{1}{2}\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})$.
We may call $\mathfrak{o}_{\mathscr{X}}$ the $\mathscr{X}$-twisted Beauville-Voisin class on $X$. From our construction, when $\mathscr{X}\to X$ is essentially trivial, it can be easily shown that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$. One may expect that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ for any $\mathscr{X}\to X$. An evidence is that if $X$ is defined over $\bar{\mathbb{Q}}$, one can show that $\mathfrak{o}_{\mathscr{X}}$ is also defined over $\bar{\mathbb{Q}}$ (See Remark 4.8) and thus $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ by the Bloch-Beilinson conjecture.
There are some useful consequences of Theorem 1.2. For instance, one can obtain a generalization of [9, Corollary 2.6]
@@
For two rigid objects $\mathcal{E},\mathcal{F}\in\mathrm{D}^{b}(\mathscr{X})$, $c_{2}(\mathcal{E})=c_{2}(\mathcal{F})\in\operatorname{CH}_{0}(X)$ if they have the same Mukai vector.
We shall mention that for in general $\mathrm{D}^{b}(\mathscr{X})$ may not have rigid objects. Moreover, motivated from the work of [21], one can define a filtration on $\mathrm{D}^{b}(\mathscr{X})$ as follows:
$\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))=\Big{\{}\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})|~{}c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{i}(X)\Big{\}}.$ (1)
Then we have
@@
For any derived equivalence $\Psi:\mathrm{D}^{b}(\mathscr{X})\to\mathrm{D}^{b}(\mathscr{Y})$ between two twisted K3 surfaces $\mathscr{X}\to X$ and $\mathscr{Y}\to Y$, we have
$\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))\Leftrightarrow\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X})).$
In other words, the filtration (1) is preserved under derived equivalences.
See [21, Corollary 2] for the untwisted case.
@@ 1.5. BV filtration on Bridgeland moduli spaces
The proof of Theorem 1.2 and Theorem 1.4 involves a dedicated study of so-called Beauville-Voisin filtration on moduli space of stable objects in $\mathrm{D}^{b}(\mathscr{X})$. Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector and $\sigma$ be a $\mathbf{v}$-generic stability condition on $\mathrm{D}^{b}(\mathscr{X})$. The (coarse) moduli space $M=M_{\sigma}(\mathscr{X},\mathbf{v})$ of $\sigma$-stable objects in $\mathrm{D}^{b}(\mathscr{X})$ with Mukai vector $\mathbf{v}$ is a smooth projective hyper-Kähler variety. Voisin has introduced a filtration $\mathbf{S}_{\bullet}^{\mathrm{B}\mathrm{V}}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$ given by
$\mathbf{S}_{i}^{\rm BV}\operatorname{CH}_{0}(M):=\left<x\in M|~{}\dim O_{x}\geq d-i\right>$
where $O_{x}=\{x^{\prime}\in M|~{}[x]=[x^{\prime}]\in\operatorname{CH}_{0}(M)\}$. From the definition, one can see such filtration only depends on the isomorphic class of $M$. On the other hand, there is a natural filtration
$\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$
given by the restriction of the filtration $\mathbf{S}_{\bullet}(\mathrm{D}^{b}(\mathscr{X}))$ to $\operatorname{CH}_{0}(M)$, i.e.
$\mathbf{S}^{\rm SYZ}_{i}\operatorname{CH}_{0}(M)=\left<x\in M|~{}x\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))\right>$ (2)
Then Theorem 1.4 actually says that filtration $\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)$ does not depend on the moduli interpretation of $M$. A natural expectation is that
$\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)=\mathbf{S}^{\rm BV}_{\bullet}\operatorname{CH}_{0}(M),$
just as the untwisted case (cf[15, Theorem 1.1]). Another goal of this paper is to verify this for arbitrary $M$. The main technical result is
@@
Assume $\mathbf{v}^{2}=2n-2\geq 0$. Let $M^{\rm l.f.}\subseteq M$ be the open locus (if non-empty) consisting of $\sigma$-stable locally free 1-fold $\mathscr{X}$-twisted sheaves. Then
$\dim\mathcal{O}_{\mathcal{E}}\cap M^{\rm l.f.}\geq n-i\Leftrightarrow\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X})).$
In particular, $\mathbf{S}_{i}^{\operatorname{BV}}\operatorname{CH}_{0}(M)=\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(M)$ if $M=M^{\rm l.f.}$.
This is proved by studying the degenerate loci of morphisms between twisted sheaves, which generalizes the work in [26] and [15]. Our method can be also applied to the case of hyper-Kähler varieties of generalized Kummer type (see some discussions in §4.15). According to the work in [7, Theorem 3.4], the locally freeness assumption can be released once $\mathscr{X}$ is derived equivalent to an untwisted K3 surface (cf[15, Proof of Theorem 1.1]). Roughly speaking, it has been shown that in that case $M$ is isomorphic to a Bridgeland moduli space which consists of slope stable locally free sheaves. This yields
@@
If $\mathrm{D}^{b}(\mathscr{X})\cong\mathrm{D}^{b}(Y)$ for some K3 surface $Y$, then $\mathbf{S}_{i}^{\rm BV}\operatorname{CH}_{0}(M)=\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(M)$
The existence of such filtration plays an important role in the study of Bloch’s conjecture for zero cycles on hyper-Kähler varieties. An advantage of Shen-Yin-Zhao type filtration $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}$ is that it is preserved under birational transformations and the graded pieces are controlled by the first one.
@@
where $\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)$ is the $s$-th graded piece of the the filtration $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$.
Another application is concerned with the existence of algebraically coisotropic subvarieties. With the same argument in [21, Theorem 0.5], one can obtain
@@
For $0\leq i\leq\frac{\mathbf{v}^{2}+2}{2}$, there exist algebraically coisotropic subvarieties
$\iota:Z_{i}\dashrightarrow B_{i}$
where $Z_{i}\subseteq M$ has codimension $i$ and the general fibers of $\iota$ are constant cycle subvarieties of $M$ with dimension $i$.
@@ 1.10. Organization of the paper
In Section 2, we review the basic results on twisted K3 surfaces and their derived categories. We discuss the twisted Chern character and its connection to the Chern character on gerbes. In Section 3, we study filtrations on the moduli space $M_{H}(\mathscr{X},\mathbf{v})$ of Gieseker stable $\mathscr{X}$-twisted torsion sheaves. Such a moduli space admits a natural Lagrangian fibration. The key observation is that there exists a finite incidence correspondence between the smooth support loci of $M_{H}(\mathscr{X},\mathbf{v})$ for various $\mathscr{X}\to X$. Section 4 is the main part of this paper. We construct the twisted Beauville-Voisin class via using the Chern class of stable twisted locally free sheaves with maximal dimensional rational equivalent orbit. By investigating the degenerate loci of morphisms between polystable twisted locally free sheaves, we show our twisted BV class is well-defined and prove Theorem 1.2. Section 5 and Section 6 are devoted to the proof of Theorem 1.4 and Theorem 1.8.

### v3

# Filtrations on the derived category of twisted K3 surfaces

@@ Abstract.
We introduce and study the Shen-Yin-Zhao filtration on derived categories of twisted K3 surfaces. A main contribution is the construction of a twisted Beauville-Voisin class $\mathfrak{o}_{\mathscr{X}}\in\operatorname{CH}_{0}(X)$ that extends fundamental results of O’Grady and Shen-Yin-Zhao [22, 23] to twisted settings. This class enables:
  • A derived equivalence-invariant filtration $\mathbf{S}_{\bullet}(\mathrm{D}^{(1)}(\mathscr{X}))$ preserved under Fourier-Mukai transforms,
  • A birational invariant filtration $\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}$ on Bridgeland moduli spaces.
We prove $\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}$ coincides with Voisin’s filtration $\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}$ (Theorem 1.4), providing a canonical candidate for the conjectural Beauville-Voisin filtration. Applications include Bloch’s conjecture for (anti)-symplectic automorphisms and existence of algebraically coisotropic subvarieties.
@@ 1.1. Filtrations on K3 category
Let $X$ be a K3 surface and let $\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)$ be the Beauville-Voisin class defined in [4]. O’Grady [22] introduced a filtration $\mathbf{S}_{\bullet}(X)$ on $\operatorname{CH}_{0}(X)$ given by
$\mathbf{S}_{i}(X)=\bigcup_{\begin{subarray}{c}\text{effective}\\ \deg([z])=i\end{subarray}}\left\{[z]+\mathbb{Z}\cdot\mathfrak{o}_{X}\right\}$
for all effective zero-cycles $z$ of degree $i$. Denote by $\mathrm{D}^{b}(X)$ the bounded derived category of coherent sheaves on $X$. Shen-Yin-Zhao [23] used O’Grady’s filtration to construct a natural increasing filtration on $\mathrm{D}^{b}(X)$:
$\mathbf{S}_{0}(\mathrm{D}^{b}(X))\subset\mathbf{S}_{1}(\mathrm{D}^{b}(X))\subset\cdots\subset\mathbf{S}_{i}(\mathrm{D}^{b}(X))\subset\cdots$
where $\mathbf{S}_{i}(\mathrm{D}^{b}(X))$ consists of objects $\mathcal{E}\in\mathrm{D}^{b}(X)$ satisfying $c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)$. They established that for any $\mathcal{E}\in\mathrm{D}^{b}(X)$,
$c_{2}(\mathcal{E})\in\mathbf{S}_{d(\mathcal{E})}(X),\quad\text{with}\quad d(\mathcal{E})=\tfrac{1}{2}\dim\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E}).$
More generally, consider a twisted K3 surface $\mathscr{X}\to X$ in the sense of [18], and denote by $\mathrm{D}^{(1)}(\mathscr{X})$ the bounded derived category of 1-fold twisted coherent sheaves (see §2.1). The existence of an analogous filtration on $\mathrm{D}^{(1)}(\mathscr{X})$ is expected. The main result of this paper is:
@@
There exists a unique degree-1 cycle class $\mathfrak{o}_{\mathscr{X}}\in\operatorname{CH}_{0}(X)$ such that for any $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$,
$c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{d(\mathcal{E})}(X),$
where $d(\mathcal{E})=\tfrac{1}{2}\dim\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})$.
We call $\mathfrak{o}_{\mathscr{X}}$ the $\mathscr{X}$-twisted Beauville-Voisin class on $X$. When the twist $\mathscr{X}\to X$ is essentially trivial, our construction yields $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$, recovering the main result of [23]. We conjecture that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ holds for any twist $\mathscr{X}\to X$. There are supporting evidence coming from the Bloch-Beilinson conjecture: when $X$ is defined over $\bar{\mathbb{Q}}$, Remark 4.5 shows that $\mathfrak{o}_{\mathscr{X}}$ is also defined over $\bar{\mathbb{Q}}$, implying $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ by the Bloch-Beilinson conjecture.
There are several useful consequences of Theorem 1.1. First, we generalize [9, Corollary 2.6]:
@@
For any two rigid objects $\mathcal{E},\mathcal{F}\in\mathrm{D}^{(1)}(\mathscr{X})$ with identical Mukai vectors,
$c_{2}(\mathcal{E})=c_{2}(\mathcal{F})\in\operatorname{CH}_{0}(X).$
Motivated by [23], we can define a derived filtration on twisted K3 categories:
$\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))=\left\{\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})\mid c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{i}(X)\right\}.$ (1)
This filtration exhibits natural invariance under derived equivalences:
@@
Let $\Psi:\mathrm{D}^{(1)}(\mathscr{X})\xrightarrow{\sim}\mathrm{D}^{(1)}(\mathscr{Y})$ be a derived equivalence between twisted K3 surfaces $\mathscr{X}\to X$ and $\mathscr{Y}\to Y$. Then for any object $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$ and index $i$,
$\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))\iff\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{Y})).$
In particular, the filtration (1) is preserved under derived equivalences.
This extends [23, Corollary 2] to the twisted setting.
@@ 1.2. BV filtration on Bridgeland moduli spaces
The proof of Theorems 1.1 and 1.3 involves a detailed study of the Beauville-Voisin filtration on the moduli space of stable objects in $\mathrm{D}^{(1)}(\mathscr{X})$. Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector and $\sigma$ a $\mathbf{v}$-generic stability condition on $\mathrm{D}^{(1)}(\mathscr{X})$. The coarse moduli space $M=M_{\sigma}(\mathscr{X},\mathbf{v})$ of $\sigma$-stable objects in $\mathrm{D}^{(1)}(\mathscr{X})$ with Mukai vector $\mathbf{v}$ is a smooth projective hyper-Kähler variety.
Voisin introduced a filtration $\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$ defined by
$\mathbf{S}_{i}^{\mathrm{BV}}\operatorname{CH}_{0}(M):=\left\langle x\in M\mid\dim\mathcal{O}_{x}\geq d-i\right\rangle,$
where $\mathcal{O}_{x}=\{x^{\prime}\in M\mid[x]=[x^{\prime}]\in\operatorname{CH}_{0}(M)\}$. This filtration depends only on the isomorphism class of $M$.
On the other hand, there is a natural filtration
$\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$
given by restricting the filtration $\mathbf{S}_{\bullet}(\mathrm{D}^{(1)}(\mathscr{X}))$ to $\operatorname{CH}_{0}(M)$, i.e.,
$\mathbf{S}^{\mathrm{SYZ}}_{i}\operatorname{CH}_{0}(M)=\left\langle\mathcal{E}\in M\mid\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))\right\rangle.$ (2)
A natural expectation, analogous to the untwisted case ([17, Theorem 1.1]), is that
$\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}(M)=\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}(M).$
Another goal of this paper is to verify this equality for arbitrary $M$. Our main technical result is:
@@
Assume $\mathbf{v}^{2}=2n-2\geq 0$. Then for any object $\mathcal{E}\in M$,
$\dim\mathcal{O}_{\mathcal{E}}\geq n-i\quad\iff\quad\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X})).$
In particular, $\mathbf{S}_{i}^{\mathrm{BV}}\operatorname{CH}_{0}(M)=\mathbf{S}_{i}^{\mathrm{SYZ}}\operatorname{CH}_{0}(M)$.
This is proved by studying degeneracy loci of morphisms between twisted sheaves, generalizing the work in [28] and [17]. Our method also applies to hyper-Kähler varieties of generalized Kummer type (see §5.5). A consequence is:
@@
Let $Y$ be a hyper-Kähler variety of $K3^{[n]}$-type. If its algebraic Markman-Mukai lattice contains an isotropic vector, then $\mathbf{S}_{\bullet}^{\mathrm{BV}}\operatorname{CH}_{0}(Y)$ is a birational invariant, i.e. for any birational map $f\colon Y\dashrightarrow Y^{\prime}$ with $Y^{\prime}$ hyper-Kähler,
$f_{\ast}(\mathbf{S}_{\bullet}^{\mathrm{BV}}\operatorname{CH}_{0}(Y))=\mathbf{S}_{\bullet}^{\mathrm{BV}}\operatorname{CH}_{0}(Y^{\prime}).$
In particular, if $\operatorname{rank}(\operatorname{NS}(Y))\notin\{2,3\}$, then $\mathbf{S}_{\bullet}^{\mathrm{BV}}\operatorname{CH}_{0}(Y)$ is a birational invariant.
This filtration plays an important role in studying Bloch’s conjecture for zero-cycles on hyper-Kähler varieties. An advantage of the Shen-Yin-Zhao filtration $\mathbf{S}_{\bullet}^{\mathrm{SYZ}}\operatorname{CH}_{0}=\mathbf{S}_{\bullet}^{\mathrm{BV}}\operatorname{CH}_{0}$ is that its graded pieces are controlled by the first one. This yields
@@
where $\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)$ is the $s$-th graded piece of the the filtration $\mathbf{S}_{\bullet}^{\rm BV}\operatorname{CH}_{0}(Y)$.
Another application concerns algebraically coisotropic subvarieties. Adapting the same argument in [23, Theorem 0.5] gives:
@@
Let $M$ be a Bridgeland moduli space as above. For $0\leq i\leq\frac{\mathbf{v}^{2}+2}{2}$, there exist algebraically coisotropic subvarieties $\iota\colon Z_{i}\dashrightarrow B_{i}$ with $Z_{i}\subseteq M$ of codimension $i$, whose general fibers are constant cycle subvarieties of dimension $i$.
@@ 1.3. Organization of the paper
In Section 2, we establish foundational concepts for twisted K3 surfaces $\mathscr{X}\to X$, including $\mu_{m}$-gerbes, twisted sheaves of weight $d$, and derived categories $\mathrm{D}^{(d)}(\mathscr{X})$. We develop the twisted Chern character $\operatorname{ch}_{\mathscr{X}}:K_{0}(\mathscr{X})\to\mathrm{CH}^{*}(X)$ and study the properties in Lemma 2.1. The twisted Mukai lattice $\widetilde{\mathrm{H}}(\mathscr{X})=\exp(\mathbf{B})\cdot\widetilde{\mathrm{H}}(X)$ and Bridgeland stability conditions are systematically reviewed.
Section 3 analyzes cycles on moduli spaces $M_{H}(\mathscr{X},\mathbf{v})$ of Gieseker stable $\mathscr{X}$-twisted torsion sheaves with primitive Mukai vector $\mathbf{v}=(0,L,s)$. We prove the Lagrangian fibration $M_{H}(\mathscr{X},\mathbf{v})\to|L|$ admits dense constant cycle Lagrangian subvarieties (Proposition 3.1), and establish the correspondence $c_{2}(\mathcal{E})\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$ for $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$ (Proposition 3.3). Key to this is Theorem 3.5, which shows:
$c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)\iff\dim(\mathcal{O}_{\mathcal{E}}\cap M^{\mathrm{sm}}_{\mathscr{X}})\geq\mathrm{d}(\mathcal{E})-i$
via finite incidence correspondences between smooth loci of moduli spaces for different twists.
Section 4 presents our core construction: the twisted Beauville-Voisin class $\mathfrak{o}_{\mathscr{X}}\in\mathrm{CH}_{0}(X)$ defined via Chern classes of stable twisted locally free sheaves on constant cycle Lagrangian subvarieties (Definition 4.1). Using degeneracy loci of morphisms between polystable sheaves (Definition 4.9), we prove $\mathfrak{o}_{\mathscr{X}}$ is well-defined and independent of choices (Proposition 4.6). This culminates in Theorem 1.1:
$c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$
for all $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$, generalizing [23] to twisted surfaces.
Sections 5 and 6 complete the main results:
  1. -
    Theorem 1.3 (derived invariance of $\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))$ via Beauville-Voisin groups (Proposition 5.1)
  2. -
    Theorem 1.4 ($\dim\mathcal{O}_{\mathcal{E}}\geq n-i\iff\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))$ unifying SYZ and BV filtrations)
  3. -
    Theorem 1.6 (Bloch’s conjecture for (anti)-symplectic birational maps), proved using birational motives (Lemma 6.3) and the identity $\mathfrak{h}^{\circ}(Y)\cong\mathfrak{h}^{\circ}(X^{[n]})$ (Proposition 6.2).
In the appendix, we extend Bayer-Macri’s result in [3] to twisted K3 surfaces, proving that Bridgeland moduli spaces $M_{\sigma}(\mathscr{X},\mathbf{v})$ are isomorphic to moduli of Gieseker stable locally free twisted sheaves (Theorem 7.2).